OpenStudy (baldymcgee6):

Inverse function of y = e^x/(1+e^x)

OpenStudy (baldymcgee6):

\[y = e^x/(1+e^x)\]

OpenStudy (baldymcgee6):

How do I solve for x?

OpenStudy (anonymous):

Order all by normalizing in compared to e^x and you get\[e ^{x} (1-y) = y \] Than just devide by (1-y) take a logirthm and solve for x.

OpenStudy (baldymcgee6):

@Botev what do you mean by normalizing?

OpenStudy (anonymous):

\(\large y = \frac{e^x}{1+e^x} \rightarrow \frac{1}{y}=\frac{1+e^x}{e^x} \) \(\large \frac{1}{y}=\frac{1+e^x}{e^x} = \frac{1}{e^x}+\frac{e^x}{e^x}=e^{-x}+1 \) \(\large \frac{1}{y}-1=e^{-x} \rightarrow \frac{1-y}{y}=e^{-x} \) can you take it from here to solve for x?

OpenStudy (baldymcgee6):

yes, thank you dpalnc!